# FusionSORT: Fusion Methods for Online Multi-object Visual Tracking

### v1

@@ I Introduction
Data association is very challenging in multi-object visual tracking due to challenges such as miss-detections due to occlusions, appearance changes, or noisy detections. To solve the association task between the predicted tracklet bounding box and the detection bounding box, many works use either one or a combination of strong cues, motion and appearance information, since these cues provide powerful instance-level discrimination. The motion information is usually computed using intersection-over-union (IoU) [1] or Mahalanobis distance [9]. The appearance information is usually leveraged from trained deep learning models [10][11]. The faster trackers such as [1][4][12] use only the motion information for the association task. Many trackers use a combination of motion and appearance information [2][3][5][13][14][15][16], generally with better performance but with compromised speed. In addition to the strong cues, some works also incorporate weak cues such as height state, confidence state and/or velocity direction for the association task [17][18][12] to compensate for the strong cues, especially in challenging situations such as occluded and crowded scenes. The key gap that is missing in the literature is the comparative studies on the effectiveness of the strategies for fusing different cues such as strong cues and/or weak cues. Different trackers use different fusion methods; however, there is no work in the literature, to the best of our knowledge, which investigates and compares the different fusion methods for data association in multi-object visual tracking.
@@
    We investigate four different fusion methods for associating detections to tracklets in multi-object visual tracking, including minimum, weighted sum based on IoU, KF gating, and hadamard product of costs.
@@ II-B Data Association
Data association in multi-object visual tracking is highly challenging due to factors like missed detections caused by occlusions, changes in object appearance, and noisy input detections. Identifying the temporally stable properties of objects is crucial to effectively associate predicted tracklet bounding boxes to detection bounding boxes in video frames. Strong cues such as motion and appearance information provide powerful instance-level discrimination. The motion information is usually computed using IoU [1] or Mahalanobis distance [9]. The appearance information is usually leveraged from trained deep learning models [10][11]. Specifically, deep appearance features are extracted from image patches determined by object detection boxes using an additional deep neural network in separate appearance-based trackers [2][3][4][5]. Appearance models can also be trained jointly with object detectors in joint trackers [14][15][23][24]. For the association task, cosine distance of the extracted deep appearance features is computed as appearance distance. Weak cues such as height state, confidence state and/or velocity direction can provide informative clues that help to compensate for the discrimination of strong cues such as motion and appearance information for associating predicted tracklet boxes to new detection boxes [17][18][12].
Combining these different sources of information for associating predicted tracket boxes to new detection boxes is very crucial. Though several trackers use only motion information [1][4][12] for data association with interesting performance, the performance can be improved by fusing different cues. Minimum fusion method is used in [5], where the minimum in each element of the cost matrices of motion and appearance is used to match tracklets to new detections. Several works [16][17][25] use weighted sum, with some variations, of motion and appearance information, in which motion information is computed using IoU. Another fusion method, which we call KF gating, is also based on weighted sum of cues and is used in [14][15][3]. However, the Mahalanobis distance is used in this fusion method instead of the IoU, and is subjected to KF gating. Hadamard product is also used in [21][18] which computes element-wise multiplication of different costs. Hence, different trackers use different fusion methods to fuse different costs for associating tracklets to current detections, which is then solved by Hungarian algorithm [26] as bipartite graph matching. However, there is no work in the literature which thoroughly investigates these different fusion methods. In this work, we investigate four widely used fusion methods and demonstrate that the choice of a fusion method is key for data association in multi-object visual tracking.
@@ III Method
For state vector representation, we extend the widely used standard KF in BoT-SORT [5] with two additional states: the tracklet confidence (score) $c$ and its velocity component $\dot{c}$, following the state vector derivation approach in [28]. Accordigly, the state vector is represented as in (1)
@@
where $(x_{c},y_{c})$ denote object tracklet box’s center, while $w$, $h$ and $c$ represent the object tracklet box’s width, height, and tracklet confidence, respectively. The velocity components are denoted by $\dot{x_{c}}$, $\dot{y_{c}}$, $\dot{w}$, $\dot{h}$ and $\dot{c}$.
@@
Following the works in [5][2], we choose the noise factors as $\sigma_{p}=0.05$, $\sigma_{v}=0.00625$, and $\sigma_{m}=0.05$, since our frame rate is also 30 fps. It is worth noting that we modified $\mathbf{Q}_{k}$ and $\mathbf{R}_{k}$ according to our slightly modified state vector $\mathbf{x}_{k}$ and measurement vector $\mathbf{z}_{k}$, respectively.
@@ III-A Strong Cues
Given a probability distribution $f$ on $\mathcal{R}^{N}$, with mean $\mathbf{\mu}=(\mu_{1},\mu_{2},\mu_{3},\dots,\mu_{N})^{\mathsf{T}}$ and positive semi-definite covariance matrix $S$, the Mahalanobis distance of a point $\mathbf{z}=(z_{1},z_{2},z_{3},\dots,z_{N})^{\mathsf{T}}$ from $f$ is given as in (6)
@@
where $\mathbf{\mu}$ and $\mathbf{z}$ corresponds to our estimated tracklet mean and measurement (detection) box center positions, respectively, while $f$ corresponds to a Gaussian predicted state distribution. $S^{-1}$ denotes inverse of the covariance matrix $S$.
@@ III-C1 Minimum
where $C_{i,j}$ denotes the (i, j) element of the total cost matrix $C$. $d^{cos}_{i,j}$, $d^{iou}_{i,j}$, $d^{hiou}_{i,j}$ and $d^{conf}_{i,j}$ are the cosine distance, IoU distance, height-IoU distance and confidence distance between i-th tracklet and j-th detection, respectively. $\wedge$ denotes logical ’and’. The minimum fusion is represented by min in (13). Note that cosine distance is 1 minus cosine similarity, and the other distances generally follow similar manner, and they are explicitly formulated as in (14)(15)(16)(17).
@@
where $\|*\|_{2}$ is the 2-norm of its argument *, and . represents the dot product of tracklet averaged appearance embedding $e^{k}_{i}$ and detection appearance embedding $f^{k}_{j}$.
@@
$d^{iou}_{i,j}$ is IoU distance, also called Jaccard distance.
@@
where |||*|| ∗ | is the absolute value of *. $c^{trk}_{i}$ and $c^{det}_{j}$ denote tracklet confidence and detection score, respectively.
@@ III-C3 KF Gating
Given the appearance cost $C_{a}$, motion cost $C_{m}$, height-IoU cost $C_{h}$ and confidence cost $C_{c}$, this method also uses weighted sum for calculating the total cost matrix $C$ as in (19). However, the motion cost is computed using Mahalanobis distance, rather than IoU, and it is subjected to KF gating [8] where the gating distance is computed between KF predicted state distribution and measurements (detections). A suitable Mahalanobis distance threshold can be obtained from a table for the 0.95 quantile of the chi-square distribution with N degrees of freedom. In our method, the chi-square distribution has 2 degrees of freedom since the distance computation is done with respect to the bounding box center position only i.e. ($x_{c},y_{c}$).
@@
Even though the Mahalanobis distance is used for the first association step where fusion of different costs takes place, the IoU is used for the second association step for all the fusion methods including this fusion method to make a fair comparison. In our experiments, the IoU gives better overall performance than the Mahalanobis distance for the second association step, as shown in Table IV. Note that we did not apply any thresholding and 0.5 multiplication as in (10)(11)(12) in this fusion method.
@@ IV-A3 Implementation Details
Our visual tracking algorithm is implemented using Python and PyTorch deep learning framework, and run on Laptop with Intel(R) Core(TM) i7-10850H @ 2.70GHz, 16 GB RAM and NVIDIA GeForce RTX 2070 GPU. We use the publicly available YOLOX-X detector [38], trained by [4] for MOT17 and MOT20 datasets, and trained by [17] for DanceTrack dataset. For feature extraction, we used the publicly available models trained by [5] for MOT17 and MOT20 datasets and trained by [17] for DanceTrack dataset, based on FastReID library [10]. We use the same tracker parameters throughout our experimental analysis. Unless otherwise specified, we set high detection score threshold $\tau_{1}$ to 0.6, which is used to separate high-confident detections from low-confident detections for the first association step. We set low detection score threshold $\tau_{2}$ to 0.1 for the second association step. Detections with score lower than $\tau_{2}$ are discarded. In the linear assignment step, the matching is considered only if the detection and the tracklet similarity is not smaller than 0.8 for the first association step and is not smaller than 0.5 for the second association step. The detection score needs to be at least 0.7 to be considered for track initialization. The lost tracklets are kept for 30 frames in case they appear again before they get deleted. Note that we do not output the boxes and identities of lost tracks, as in [4][5]. In order to precisely investigate the effectiveness of the fusion methods, we do not apply any tracklet interpolation as a post-processing in our experiments.
@@ V Conclusion
In this paper, we investigate four different fusion methods for associating detections to tracklets in multi-object visual tracking by considering strong cues such as motion and appearance information as well as weak cues such as height intersection-over-union (height-IoU) and tracklet confidence information. The fusion methods include minimum, weighted sum based on IoU, Kalman filter (KF) gating, and hadamard product of costs due to the different cues. For computing confident cost, we elegantly incorporate tracklet confidence state into the state vector representation of the KF. Through extensive experiments on validation sets of MOT17, MOT20 and DanceTrack datasets, we find out that the different fusion methods have their own pros and cons. The minimum fusion method works reasonably well when fusing motion and appearance information; however, its performance decreases when incorporating weak cues. The incorporation of weak cues also degrades the tracking performance when using the hadamard fusion method. The tracking performance increases when using weak cues along with strong cues when using weighted sum and KF gating fusion methods though the performance of the KF gating fusion method is generally lower on DanceTrack dataset. Hence, the weighted sum based on IoU is more favourable when using weak cues along with the strong cues. We hope that this investigative work helps the computer vision research community to use the right fusion method with given cues for data association in multi-object visual tracking.

### v3

@@ I Introduction
Data association is very challenging in multi-object visual tracking due to challenges such as miss-detections due to occlusions, appearance changes, or noisy detections. To solve the association task between the predicted tracklet bounding box and the detection bounding box, many works use either one or a combination of strong cues, motion and appearance information, since these cues provide powerful instance-level discrimination. The motion information is usually computed using intersection-over-union (IoU) and its variants [9] or Mahalanobis distance [10]. Normalized Euclidean distance between bounding box centers of detection and tracklet is also used in [7]. The appearance information is usually leveraged from trained deep learning models [11][12]. The faster trackers such as [1][4][13] use only the motion information for the association task. Many trackers use a combination of motion and appearance information [2][3][5][14][15][16][17], generally with better performance but with compromised speed. In addition to the strong cues, some works also incorporate weak cues such as height state, confidence state and/or velocity direction for the association task [18][19][13] to compensate for the strong cues, especially in challenging situations such as occluded and crowded scenes. The key gap that is missing in the literature is the comparative study on the effectiveness of the strategies for fusing different cues such as strong cues and/or weak cues. Different trackers use different fusion methods; however, there is no work in the literature, to the best of our knowledge, which investigates and compares the different fusion methods for data association in multi-object visual tracking.
@@
    We investigate four widely used different fusion methods for associating detections to tracklets in multi-object visual tracking, including minimum, weighted sum based on IoU, KF gating, and hadamard product of costs.
@@ II-B Data Association
Data association in multi-object visual tracking is highly challenging due to factors like missed detections caused by occlusions, changes in object appearance, and noisy input detections. Identifying the temporally stable properties of objects is crucial to effectively associate predicted tracklet bounding boxes to detection bounding boxes in video frames. Strong cues such as motion and appearance information provide powerful instance-level discrimination. The motion information is usually computed using IoU and its variants [9] or Mahalanobis distance [10]. Normalized Euclidean distance between bounding box centers of detection and tracklet is also used in [7]. The appearance information is usually leveraged from trained deep learning models [11][12]. Specifically, deep appearance features are extracted from image patches determined by object detection boxes using an additional deep neural network in separate appearance-based trackers [2][3][4][5]. Appearance models can also be trained jointly with object detectors in joint trackers [15][16][24][25]. For the association task, cosine distance of the extracted deep appearance features is computed as appearance distance. Weak cues such as height state, confidence state and/or velocity direction can provide informative clues that help to compensate for the discrimination of strong cues such as motion and appearance information for associating predicted tracklet boxes to new detection boxes [18][19][13].
Combining these different sources of information for associating predicted tracket boxes to new detection boxes is very crucial. Though several trackers use only motion information [1][4][13] for data association with interesting performance, the performance can be improved by fusing different cues. Minimum fusion method is used in [5], where the minimum in each element of the cost matrices of motion and appearance is used to match tracklets to new detections. Several works [17][18][26] use weighted sum, with some variations, of motion and appearance information, in which motion information is computed using IoU. Another fusion method, which we call KF gating, is also based on weighted sum of cues and is used in [15][16][3]. However, the Mahalanobis distance is used in this fusion method instead of the IoU, and is subjected to KF gating. Hadamard product is also used in [22][19] which computes element-wise multiplication of different costs. Hence, different trackers use different fusion methods to fuse different costs for associating tracklets to current detections, which is then solved by Hungarian algorithm [27] as bipartite graph matching. However, there is no work in the literature which thoroughly investigates these different fusion methods. In this work, we investigate four commonly used fusion methods and demonstrate that the choice of a fusion method is key for data association in multi-object visual tracking.
@@ III Method
For state vector representation, we extend the widely used standard KF in [5] with two additional states: the tracklet confidence (score) $c$ and its velocity component $\dot{c}$, following the state vector derivation approach in [29]. Accordigly, the state vector is represented as in (1)
@@
where $(x_{c},y_{c})$ denote object tracklet box’s center, while $w$, $h$ and $c$ represent the object tracklet box’s width, height, and tracklet confidence, respectively. The velocity components are denoted by $\dot{x_{c}}$, $\dot{y_{c}}$, $\dot{w}$, $\dot{h}$ and $\dot{c}$. Tracklet box size can change dramatically when predicting inactive (lost) tracks for multiple frames without state update, which hinders re-activation after occlusion. To overcome this, we apply height preservation, setting the derivative $\dot{h}$ to zero before the KF prediction step, similar to [9][5][4]. Similarly, we also apply width preservation ($\dot{w}=0$) and tracklet confidence preservation ($\dot{c}=0$) before the KF prediction step in our experiments. Note that the tracklet confidence preservation has less impact on some sequences when compared to the others.
@@
Following the works in [5][2], we choose the noise factors as $\sigma_{p}=0.05$, $\sigma_{v}=0.00625$, and $\sigma_{m}=0.05$, since our frame rate is also 30 fps. It is worth noting that we modified $\mathbf{Q}_{k}$ and $\mathbf{R}_{k}$ according to our slightly modified state vector $\mathbf{x}_{k}$ and measurement vector $\mathbf{z}_{k}$, respectively. Noise Scale Adaptive (NSA) KF [3], $\mathbf{R}_{NSA}=(1-z_{c})\mathbf{R}$, did not help in our experiments. Ideally, the higher the detection confidence, the smaller the adapted measurement noise covariance $\mathbf{R}_{NSA}$ and the more influence has the detection on the track state update.
@@ III-A Strong Cues
It is possible to incorporate the uncertainty of the motion estimation into the distance measure since the KF is used as a motion model. Given a probability distribution $f$ on $\mathcal{R}^{N}$, with mean $\mathbf{\mu}=(\mu_{1},\mu_{2},\mu_{3},\dots,\mu_{N})^{\mathsf{T}}$ and positive semi-definite covariance matrix $S$, the Mahalanobis distance of a point $\mathbf{z}=(z_{1},z_{2},z_{3},\dots,z_{N})^{\mathsf{T}}$ from $f$ is generally given as in (6)
@@
where $\mathbf{z}$ and $\mathbf{\mu}$ correspond to measurement (detection) box center position and the projection of the estimated tracklet mean into measurement space, respectively, while $f$ corresponds to a Gaussian predicted state distribution. $S^{-1}$ denotes inverse of the projected tracklet state covariance matrix $S$ into the measurement space. Hence, ($\mathbf{\mu}$, $S$) corresponds to a track state projected into the measurement space. We use the squared Mahalanobis distance $d^{2}_{M}(\mathbf{z},f)$ in our experiments.
@@ III-C1 Minimum
where $C_{i,j}$ denotes the (i, j) element of the total cost matrix $C$. $d^{cos}_{i,j}$, $d^{iou}_{i,j}$, $d^{hiou}_{i,j}$ and $d^{conf}_{i,j}$ are the cosine distance, IoU distance, height-IoU distance and confidence distance between i-th tracklet and j-th detection, respectively. $\wedge$ denotes logical ’and’. The minimum fusion is represented by min in (13). Note that cosine distance is 1 minus cosine similarity, and the other distances generally follow similar manner, and they are explicitly formulated as in (14), (15), (16), and (17).
@@
where $\|*\|_{2}$ is the 2-norm of its argument *, and . represents the dot product of tracklet averaged appearance embedding $e^{k}_{i}$ and detection appearance embedding $f^{k}_{j}$. Note that $d^{cos}_{i,j}\in[0,2]$.
@@
$d^{iou}_{i,j}\in[0,1]$ is IoU distance, also called Jaccard distance.
@@
where |||*|| ∗ | is the absolute value of *. $c^{trk}_{i}$ and $c^{det}_{j}$ denote tracklet confidence and detection score, respectively. Note that $d^{hiou}_{i,j}\in[0,1]$ and $d^{conf}_{i,j}\in[0,1]$.
@@ III-C3 KF Gating
Given the appearance cost $C_{a}$, motion cost $C_{m}$, height-IoU cost $C_{h}$ and confidence cost $C_{c}$, this method also uses weighted sum for calculating the total cost matrix $C$ as in (19). However, the motion cost is computed using Mahalanobis distance, rather than IoU, and it is subjected to KF gating [8] (preventing unlikely assignments) where the gating distance is computed between KF predicted state distribution and measurements (detections). A suitable Mahalanobis distance threshold can be obtained from a table for the 0.95 quantile of the chi-square distribution with N degrees of freedom. In our method, the chi-square distribution has 2 degrees of freedom since the distance computation is done with respect to the bounding box center position only i.e. ($x_{c},y_{c}$).
@@
Even though the Mahalanobis distance is used for the first association step where fusion of different costs takes place, the IoU is used for the second association step for all the fusion methods including this fusion method to make a fair comparison. In our experiments, the IoU gives better overall performance than the Mahalanobis distance for the second association step, as shown in Table IV. Note that we did not apply any thresholding and 0.5 multiplication as in (10), (11), and (12) in this fusion method.
@@ IV-A3 Implementation Details
Our visual tracking algorithm is implemented using Python and PyTorch deep learning framework, and run on Laptop with Intel(R) Core(TM) i7-10850H @ 2.70GHz, 16 GB RAM and NVIDIA GeForce RTX 2070 GPU. We use the publicly available YOLOX-X detector [39], trained by [4] for MOT17 and MOT20 datasets, and trained by [18] for DanceTrack dataset. For feature extraction, we used the publicly available models trained by [5] for MOT17 and MOT20 datasets and trained by [18] for DanceTrack dataset, based on FastReID library [11]. We use the same tracker parameters throughout our experimental analysis, which were mostly set empirically. Unless otherwise specified, we set high detection score threshold $\tau_{1}$ to 0.6, which is used to separate high-confident detections from low-confident detections for the first association step. We set low detection score threshold $\tau_{2}$ to 0.1 for the second association step. Detections with score lower than $\tau_{2}$ are discarded. In the linear assignment step, the matching is rejected if the detection and the tracklet similarity is smaller than 0.2 for the first association step and smaller than 0.5 for the second association step. The detection score needs to be at least 0.7 to be considered for track initialization. The lost tracks (inactive tracks) are kept for 30 frames in case they appear again before they get deleted. Note that we do not output the boxes and identities of lost tracks, as in [4][5]. In order to precisely investigate the effectiveness of the fusion methods, we do not apply any tracklet interpolation as a post-processing in our experiments.
@@ V Conclusion
In this paper, we investigate four commonly used different fusion methods for associating detections to tracklets in multi-object visual tracking by considering strong cues such as motion and appearance information as well as weak cues such as height intersection-over-union (height-IoU) and tracklet confidence information. The fusion methods include minimum, weighted sum based on IoU, Kalman filter (KF) gating, and hadamard product of costs due to the different cues. For computing confident cost, we elegantly incorporate tracklet confidence state into the state vector representation of the KF. Through extensive experiments on validation sets of MOT17, MOT20 and DanceTrack datasets, we find out that the different fusion methods have their own pros and cons. The minimum fusion method works reasonably well when fusing motion and appearance information; however, its performance decreases when incorporating weak cues. The incorporation of weak cues also degrades the tracking performance when using the hadamard fusion method. The tracking performance increases when using weak cues along with strong cues when using weighted sum and KF gating fusion methods though the performance of the KF gating fusion method is generally lower on DanceTrack dataset. Hence, the weighted sum based on IoU is more favourable when using weak cues along with the strong cues. We hope that this investigative work helps the computer vision research community to use the right fusion method with given cues for data association in multi-object visual tracking.